Sunland Inc. issues $3,500,000 of 7% bonds due in 15 years with interest payable at year-end. The current market rate of interest for bonds of similar risk is 12%. Click here to view factor tables What amount will Sunland receive when it issues the bonds? (Round factor values to 5 decimal places, e.g. 1.25124 and final answers to 0 decimal places, e.g. 458,581.) Amount received by Sunland when bonds were issued $enter a dollar amount rounded to 0 decimal places
Answer #1
Amount PV factor 12% Present value
Annual Interest 245000 6.81086 1668661
Principal 3500000 0.18270 639450
Total 2308111
Amount received by Sunland when bonds were issued = $2308111
Workings:
Annual Interest 245000 =3500000*7%
PV factor 12%:
Annual Interest 6.81086 =(1-(1.12)^-15)/0.12
Principal 0.18270 =1/1.12^15